# appendixReferences

### v1

@@ D.3 Latent Concepts
As mentioned in the main text, concept learning is an important component of RAISE. This section shows the interpolation results of latent concepts on all image configurations and the correspondences between latent concepts and real attributes in Figures 9 and 10. In most cases, RAISE learns independent latent concepts and the binary matrix $\bm{M}$ that can accurately reflect the concept-attribute correspondences. RAISE does not assign the latent concepts encoding object rotations to any attribute since the noise attributes are not included in rule annotations. This experiment illustrates the interpretability of the latent concepts, which benefits the prediction of correct answers and the following experiment of odd-one-out.
@@ D.4 Odd-one-out in RPM
In this experiment, we give the additional results of odd-one-out on different configurations where RAISE picks out odd images interpretably via prediction deviations of concepts. Figure 11 visualizes the experimental results of the odd-one-out problems. We observe that RAISE will display larger deviations at odd concepts, which is important evidence when solving odd-one-out problems. It should be pointed out that forming such concept-level prediction deviations requires the model to parse independent concepts correctly. RAISE can apply the atomic rules in the global knowledge set to tasks like out-one-out and has interpretability in generative abstract reasoning.
@@ D.5 Strategy of Answer Selection
Table 6: The accuracy (%) of RAISE using different answer selection strategies.
@@
In this experiment, we evaluate RAISE with two strategies of answer selection: comparing candidates and predictions in pixel space (RAISE-pixel) and latent space (RAISE-latent). The experimental results are given in Table 6, which reports higher accuracy when candidates and predictions are compared in latent space. Due to the noise in attributes, there can be multiple solutions to a generative RPM problem. Let the answer of an RPM be the image having two triangles. By generating two triangles in various positions, the answer images may significantly differ from each other in the pixel space. However, they still point to the same concepts Number=2 and Shape=Triangle in the latent space. Therefore, RAISE selects answers by comparing candidates and predictions in the latent space rather than the pixel space.
@@
Figure 9: Concept learning on RAVEN. The table shows the interpolation results of latent concepts and the binary matrices learned in the ELBO to indicate the correspondence between concepts and attributes on L-R and U-D.
Figure 10: Concept learning on RAVEN. The table shows the interpolation results of latent concepts and the binary matrices learned in the ELBO to indicate the correspondence between concepts and attributes on O-IG, 2$\times$2Grid, and 3$\times$3Grid.
Figure 11: Odd-one-out based on RPMs. The plots display how to construct odd-one-out tests from different configurations of RPMs and how to find the odd image from the prediction deviations in concepts.
[…]

### v3

@@ D.3 Latent Concepts
As mentioned in the main text, concept learning is an important component of RAISE. This section shows the interpolation results of latent concepts on all image configurations and the correspondences between latent concepts and real attributes in Figures 9 and 10. In most configurations, RAISE can learn independent latent concepts and the binary matrix $\bm{M}$ that accurately reflects the concept-attribute correspondences. RAISE does not assign the latent concepts encoding object rotations to any attribute since the noise attributes are not included in rule annotations. This experiment illustrates the interpretability of the acquired latent concepts, which benefits the prediction of correct answers and the following experiment of odd-one-out.
@@ D.4 Odd-one-out in RPM
In this experiment, we provide the additional results of odd-one-out on different configurations where RAISE picks out rule-breaking images interpretably via prediction errors on latent concepts. Figure 11 visualizes the experimental results of odd-one-out. RAISE will display larger prediction errors at odd concepts, which is important evidence when solving odd-one-out problems. It should be pointed out that forming such concept-level prediction errors requires the model to parse independent latent concepts and conduct concept-specific abstract reasoning correctly. RAISE can apply the atomic rules in the global knowledge set to tasks like out-one-out and has interpretability in generative abstract reasoning.
@@ D.5 Strategy of Answer Selection
Table 7: The accuracy (%) using different strategies of answer selection.
@@
In this experiment, we evaluate RAISE with two strategies of answer selection: comparing candidates and predictions in pixel space (RAISE-pixel) and latent space (RAISE-latent). Table 7 reports higher accuracy when candidates and predictions are compared in latent space. Due to the noise in attributes, there can be multiple solutions to a generative RPM problem. Assume that the answer to an RPM is the image having two triangles, the answer images may significantly differ from each other in the pixel space by generating two triangles in various positions. However, they still point to the same concepts Number=2 and Shape=Triangle in the latent space. Therefore, selecting answers by comparing candidates and predictions in the latent space can be more accurate than comparing in the pixel space.
Appendix E Discussion on Bayesian and Neural Concept Learning
[…]
Solving multi-solution problems. There can be multiple solutions for one generative reasoning problem due to the noise in data. DLVMs can handle multi-solution problems by stochastic sampling from the generative and inference processes. For example, RAISE can produce results different from the original sample but still follow the correct rules. Instead of making deterministic predictions, DLVMs attempt to provide probabilities of generating specific answers and capture randomness and uncertainty in abstract reasoning.
[…]
@@ Appendix E Discussion on Bayesian and Neural Concept Learning
Figure 9: Concept learning on RAVEN. The table shows the interpolation results of latent concepts and the binary matrices indicating the correspondence between concepts and attributes on L-R and U-D.
Figure 10: Concept learning on RAVEN. The table shows the interpolation results of latent concepts and the binary matrices indicating the correspondence between concepts and attributes on O-IG, 2$\times$2Grid, and 3$\times$3Grid.
Figure 11: Odd-one-out based on RPMs. The plots display how to construct odd-one-out tests from different configurations of RPMs and how to find the odd image according to the prediction errors on latent concepts.